OpenStudy (anonymous):

he perimeter of a triangle is 61 centimeters. the longest side is 3 centimeters(s) less than the sum of the other two sides.Twice the shortest sides is 19 centimeters less than the longest side.find the lengths of each side.

OpenStudy (anonymous):

Let s, L and b be the shortest side, the longest side and the remaining side respectively. Solve the following for s, b and L:\[\{s+b+L=61,L=s+b-3,2s=L-19\} \]\[\{s\to 5,b\to 27,L\to 29\} \]